a group of 100 people have the blood types A20 B24 AB10 O46. If 1 person is selected at random what is the probability that the person has type AB blood : the person has either thype A of type O blood : the person does not have type B blood : the person has neither thype B nor type O blood ?
Answer by stanbon(75887)

a group of 100 people have the blood types A20 B24 AB10 O46.
If 1 person is selected at random
what is the probability
that the person has type AB blood = 10/100 = 1/10
----------
the person has either type A of type O blood = (20/100)+(46/100) = 33/50
----------
: the person does not have type B blood = 1-(24/100) = 76/100 = 19/25
---------
: the person has neither type B nor type O blood = (24/100)(46/100)=0.1104
==============
Cheers,
Stan H.